OpenStudy (anonymous):

Evaluate x+y/5 for x=15and y=10

hero (hero):

answer should be 5

hero (hero):

(15+10)/5 = 25/5 = 5

OpenStudy (anonymous):

the answers is 5

hero (hero):

Cool beans

OpenStudy (anonymous):

cool bean

hero (hero):